OpenStudy (tanabugg):

What the surface area of a cone with a diameter of 8 and height of 7

OpenStudy (tanabugg):

@rebeccaxhawaii

OpenStudy (mustbethestars):

Okay first we need to analyze the formula for S.A. of a cone. Do you know the formula?

OpenStudy (tanabugg):

Yes

OpenStudy (mustbethestars):

Okay what is it :3 ?

OpenStudy (tanabugg):

I'm gonna put it in the equation thing

OpenStudy (mustbethestars):

Okie.

OpenStudy (tanabugg):

\[\pi \times r( r + \sqrt{h ^{2 }}\]

OpenStudy (tanabugg):

+ radius2

OpenStudy (tanabugg):

It wouldn't let me do that part at the end idk

OpenStudy (mustbethestars):

Correct now cut the diameter in half to get your radius so 8/2 = 4

rebeccaxhawaii (rebeccaxhawaii):

http://prntscr.com/a9zksz

OpenStudy (tanabugg):

right

rebeccaxhawaii (rebeccaxhawaii):

plug it in so your radius is 4 and pi is 3.14

OpenStudy (mustbethestars):

Smh

rebeccaxhawaii (rebeccaxhawaii):